A travelling wave in a string is given by y=0.03sin (2.2x-3.5t) where y and x are in meters and t is in seconds find the amplitude,wavelength, frequency , speed of the wave


1
Expert's answer
2021-09-14T11:02:30-0400

The equation of a travelling wave in general form is the following


"y = A\sin \left(\dfrac{2\pi}{\lambda}x-2\pi f t \right)"

where "A" is the amplitude, "\lambda" is the wavelength, "f" is the frequency, and the speed of the wave is given as follows:


"v = \lambda f"

Comparing the general expression with the given one, find the paremeters. Amplitude: "A = 0.03m", wavelenght: "\lambda = \dfrac{2\pi}{2.2}m\approx 2.86m", frequency: "f = \dfrac{3.5}{2\pi}Hz \approx 0.557Hz", speed: "v =2.86m\cdot 0.557Hz \approx 1.59m/s".


Answer. Amplitude: "A = 0.03m", wavelenght: "\lambda = 2.86m", frequency: "f = 0.557Hz", speed: "v = 1.59m\/s".